Tara pushes a 10.5 kg block of ice 11.0 m across the floor with a force of 235 N. What is the ice block’s velocity the moment she stops pushing it?
1
Expert's answer
2020-11-16T07:48:33-0500
"K=W\\0.5mv^2=Fs\\0.5(10.5)v^2=(235)(11)\\v=22.2\frac{m}{s}"